OpenStudy (anonymous):

Solve for n and select the correct answer. (15n+3)/2=6 n = 14 n = -1 n = -3 n = 19

OpenStudy (cggurumanjunath):

cross multiply !

OpenStudy (cggurumanjunath):

(15n+3)=6*2=12

OpenStudy (cggurumanjunath):

now can u solve for n ? and tell me the final answer .

OpenStudy (anonymous):

i'm confused

OpenStudy (cggurumanjunath):

(15n+3)=6*2=12 this step is clear right ? @ilyautumn

OpenStudy (anonymous):

i guess kinda. Why are there 2 equal signs?

OpenStudy (cggurumanjunath):

6*2 =12 right !

OpenStudy (anonymous):

yes

OpenStudy (cggurumanjunath):

(15n+3)=12

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

then subtract 3

OpenStudy (anonymous):

from both sides?

OpenStudy (cggurumanjunath):

yes, now can u solve for n ? and tell me the final answer .

OpenStudy (cggurumanjunath):

15n+3=12 15n+3-3=12-3

OpenStudy (cggurumanjunath):

then

OpenStudy (anonymous):

15n=9

OpenStudy (cggurumanjunath):

rhs should be -6 check the q properly !

OpenStudy (cggurumanjunath):

Solve for n and select the correct answer. ((15n+3)/2=-6

OpenStudy (cggurumanjunath):

not Solve for n and select the correct answer. (15n+3)/2=6

OpenStudy (anonymous):

what????

OpenStudy (cggurumanjunath):

rhs = right hand side of the given equation. -6 ; not +6

OpenStudy (anonymous):

-1

OpenStudy (cggurumanjunath):

is rhs = -6 then n=-1.

OpenStudy (cggurumanjunath):

rhs is - 6 right ?